# A delayed interior area-to-height estimate for the Curve Shortening Flow

### v1

@@ 1.2 Main results
Let $\nu=u_{0}\mathscr{L}^{1}+\nu_{\mathrm{sing}}\in\mathcal{M}_{*}(\mathbb{R})$ be a non-atomic real-valued Radon measure, decomposed into its absolutely continuous $\nu_{\mathrm{ac}}=u_{0}\mathscr{L}^{1}$ and its singular $\nu_{\mathrm{sing}}$ part; here $\mathscr{L}^{1}$ is the Lebesgue measure on $\mathbb{R}$. Let $\Omega=\mathbb{R}\setminus\operatorname{supp}(\nu_{\mathrm{sing}})$ and suppose that $u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\Omega)$ for some $1\leq p<\infty$ (recall that the weakest case of $p=1$ always holds by definition). Then there exists a Graphical Curve Shortening Flow $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ which observes the following modes of convergence as $t\searrow 0$:
[…]
@@
    (1.16) $u(t)\to u_{0}\in\mathrm{L}^{p}(\mathbb{R}).$
@@ 1.3 Acknowledgements
The author would like to thank his supervisor Prof. Peter M. Topping for the countless and invaluable discussions and suggestions which lead to this paper, and to thank Drs. Ben S. Lambert and Huy T. Nguyễn for a number of helpful comments.
@@ 2 A local Harnack quantity
Let $u_{0}\in\mathrm{C}^{0,1}(J_{1})$ be a Lipschitz function; we construct the Lipschitz curve $\gamma_{0}:\mathbb{R}\mapsto\mathbb{R}^{2}$ as follows (see Fig. 2.1): For $x\in J_{1}$, set $\gamma_{0}(x)=(x,u_{0}(x)\!)$; for $x=1+y\in[1,\infty)$, set $\gamma_{0}(x)=(1,u_{0}(1)+y)$; and for $x=-1-y\in(-\infty,-1]$, set $\gamma_{0}(x)=(-1,u_{0}(-1)+y)$. Then the Chou–Zhu flow $[0,\infty)\ni t\mapsto\gamma(t)\subset\mathbb{R}^{2}$ of $\gamma_{0}=:\gamma(0)$ admits a graphical representation for some $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,\infty)\!)$ with $u(t)\to u_{0}\in\mathrm{C}^{0}_{\mathrm{loc}}(J_{1})$ as $t\searrow 0$. Moreover, $u(y,t),\pm\partial_{x}u(y,t)\to\infty$ locally uniformly in $t\in(0,\infty)$ as $y\to\pm 1$.
@@
Figure 2.1: Construction of $\gamma_{0}:\mathbb{R}\mapsto\mathbb{R}^{2}$ from a given $u_{0}\in\mathrm{C}^{0,1}(J_{1};\mathbb{R}_{\geq 0})$.
Proof.
@@
  1. (1)
    First, we show that each $\gamma(t)$ for $t\in[0,T]$ is asymptotic to the vertical lines $V_{\pm 1}=\left\{\pm 1\right\}\times\mathbb{R}\subset\mathbb{R}^{2}$. Applying the Avoidance Principle (Theorem A.7) to $\gamma$ and to huge shrinking circles placed either to the left of $V_{-1}$ or to the right of $V_{1}$ shows that $\gamma(t)$ must live within the strip $\overline{J_{1}}\times\mathbb{R}\subset\mathbb{R}^{2}$. Applying the Avoidance Principle again with huge circles placed below the $x$–axis shows that $\gamma(t)$ must actually live within the half-strip $\overline{J_{1}}\times\mathbb{R}_{\geq 0}$. The Strong Maximum Principle implies that $\gamma(t)$ cannot touch the boundary of this half-strip.
@@
    (2.2) $\mathfrak{o}(s)=\left\{(x,y)\in J_{1}\times\mathbb{R}\,\middle|\,\cosh\left(\frac{\pi y}{2}\right)=e^{-\pi^{2}s/4}\cos\left(\frac{\pi x}{2}\right)\right\}\subset\mathbb{R}^{2}.$
    Let $L>\|u_{0}\|_{\mathrm{L}^{\infty}(J_{1})}$; by translating $\mathfrak{o}(s)$ upwards to touch the line $H_{L}=\mathbb{R}\times\{L\}\subset\mathbb{R}^{2}$ from above, and by considering the flow $[0,|s|)\ni t\mapsto\mathfrak{o}(t+s)$, the Avoidance Principle ensures that the flow of $\gamma_{0}$ must lie outside the flow of the Angenent oval solution—see Fig. 2.2—for any $s<0$. In particular, a direct calculation (see e.g. [Sob24, §2.5.2]) shows that as $s\searrow-\infty$, the Angenent oval solutions approach the vertical translation of the Grim Reaper solution $[0,\infty)\ni t\mapsto\mathfrak{g}(t)\subset\mathbb{R}^{2}$, given by
    (2.3) $\mathfrak{g}(t)=\left\{(x,y)\in J_{1}\times\mathbb{R}\,\middle|\,e^{-\pi y/2}=e^{-\pi^{2}t/4}\cos\left(\frac{\pi x}{2}\right)\right\}\subset\mathbb{R}^{2},$
@@
    Next, we show that the curve $\gamma(t)$ is given by a graph in two senses. We first apply Angenent’s Intersection Principle (Theorem A.6) to $\gamma$ and to the static flow of the vertical lines $V_{r}=\left\{r\right\}\times\mathbb{R}\subset\mathbb{R}^{2}$ for $r\in J_{1}$; see Fig. 2.2. There is initially a single intersection coming from the graphical representation $\gamma_{0}=\operatorname{graph}(u_{0})$, and for strictly positive times the curve $\gamma(t)$ must cross $V_{r}$ at least once; thus Angenent’s Intersection Principle ensures that the curve $\gamma(t)$ cross each $V_{r}$ exactly once. Hence $\gamma(t)=\operatorname{graph}(u(t)\!)$ is a graph of a function $u(t):J_{1}\mapsto\mathbb{R}_{>0}$.
@@
    We second apply Angenent’s Intersection Principle to $\gamma$ and to the horizontal lines $H_{r^{\prime}}=\mathbb{R}\times\left\{r^{\prime}\right\}\subset\mathbb{R}^{2}$; see Fig. 2.2. Let $\underline{r}^{\prime}\gg 1$ be such that $u(0,t)<\underline{r}^{\prime}$ for all $t\in[0,T]$. Then, similarly to before, we have for all ${r^{\prime}}\geq\underline{r}^{\prime}$ that each $\gamma(t)$ for $t\in[0,T]$ intersects $H_{r^{\prime}}$ exactly twice, and therefore that the ends of $\gamma(t)$ are graphical over {±$\left\{\pm 1\right\}\times[\underline{r}^{\prime},\infty)\subset V_{\pm 1}$.
  3. (3)
@@
    Finally, we establish the convergence properties from regularity theory. We know from (2) that, say, for $t\in[0,T]$ the right-most end of $\gamma(t)$ can be written as a graph of $\widetilde{u}(t):[\underline{r}^{\prime},\infty)\mapsto\mathbb{R}$ over $\left\{1\right\}\times[\underline{r}^{\prime},\infty)\subset V_{1}$, which starts off as $\widetilde{u}(0)=0$ but which instantly lifts off from zero (i.e. $\widetilde{u}(t)>0$ for all $t\in(0,T]$). From (1), we know that the graphing function $\widetilde{u}(y,t)$ decays rapidly, say $|\widetilde{u}(y,t)|\leq\Psi(y)\to 0$, as $y\nearrow\infty$, since $\gamma(t)$ is squeezed between the vertical line $V_{1}$ and an upward-translated Grim Reaper (2.3).
@@
    Ecker & Huisken’s gradient estimate [EH91, 2.1 Theorem] implies that $\partial_{x}\widetilde{u}(y,t)$ remains uniformly bounded for, say, $(y,t)\in[\underline{r}^{\prime\prime},\infty)\times[0,T]$ for some $\underline{r}^{\prime\prime}\gg\underline{r}^{\prime}$. Standard parabolic regularity theory [^4] See, for example, [LSU88, Kry96]. For our purposes here, it is not important to clarify how the constant $C$ depends on the solution $\widetilde{u}$, as we only require the qualitative statement that $\partial_{x}\widetilde{u}(y,t)\to 0$ as $y\nearrow\infty$ locally uniformly in $t\in(0,\infty)$. In the proof of Theorem 1.13, it will be paramount to obtain quantitative estimates that depend on only weak information about the solution. Therefore, in §5.1 we will carry out a procedure similar to here, but with a little more care. then ensures that the gradient $\partial_{x}\widetilde{u}(y,t)$ of the graphing function must also decay to zero as $y\nearrow\infty$ at the same rate as $\widetilde{u}(y,t)$: $|\partial_{x}\widetilde{u}(y,t)|\leq C\Psi(y)\to 0$. Written as a graph over $J_{1}\times\{0\}$, this implies that both $u(y,t)$ and $\partial_{x}u(y,t)$ blow up to infinity as $y\nearrow 1$, locally uniformly in $t\in(0,T]$. The argument for the left-most end of $\gamma(t)$ is similar. ∎
@@
[…]
Figure 2.2: Left: Applying the Avoidance Principle (Theorem A.7) to the solid $\gamma$ and to a dashed Angenent oval solution $\mathfrak{o}$ in (1).
Right: Applying Angenent’s Intersection Principle (Theorem A.6) to the solid $\gamma$ and to a dashed vertical/horizontal line $V_{r}$/$H_{r^{\prime}}$ in (2).
@@
The proof of Lemma 2.1, in particular (1), shows the following. Let $L>\|u_{0}\|_{\mathrm{L}^{\infty}(J_{1})}$. Then the restricted solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,T])$ from Lemma 2.1 is dominated by the Grim Reaper solution (2.3) translated upwards by $L$; that is,
(2.5) $u(x,t)\leq L+\frac{\pi t}{2}-\frac{2}{\pi}\log\left(\cos\left(\frac{\pi x}{2}\right)\!\right)\leq L+\frac{\pi T}{2}-\frac{2}{\pi}\log\left(\cos\left(\frac{\pi x}{2}\right)\!\right)$
@@
Let $u_{0}\in\mathrm{C}^{0,1}(J_{1};\mathbb{R}_{\geq 0})$ be a positive Lipschitz function on $J_{1}$. The local GCSF starting from $u_{0}$ is the solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,\infty)\!)$ to the GCSF constructed in Lemma 2.1. Note that $u$ extends to a function in $\mathrm{C}^{0}_{\mathrm{loc}}(J_{1}\times[0,\infty)\!)$.
@@
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,\infty)\!)$ be a local GCSF. Define the area and angle [^5] The quantity $\phi(y,t)$ is the angle that the tangent to the curve $\gamma_{u}(t)$ at $(y,u(y,t)\!)$ makes with the downward vertical direction. functions $\mathcal{A}:J_{1}\times[0,\infty)\mapsto[0,\infty]$ and $\phi:J_{1}\times(0,\infty)\mapsto(0,\pi)$ by
(2.8) $\mathcal{A}(y,t)$ $=\int_{x\in(-1,y)}u(x,t)\,\mathrm{d}x$
@@
Figure 2.3: Illustration of $\mathcal{A}(y,t)$ and $\phi(y,t)$.
The content of the following proposition is to confirm that $\mathcal{H}$ is smooth on $J_{1}\times(0,\infty)$, and extends continuously to $\overline{J_{1}}\times[0,\infty)$ with well-controlled boundary behaviour.
@@
(2.14) $\mathcal{A}(-1,t)=0\qquad\text{ and }\qquad\mathcal{A}(1,t)=\|u(t)\|_{\mathrm{L}^{1}(J_{1})}.$
The boundedness of $\phi(y,t)$ and the locally uniform in $t\in(0,T]$–convergence of $\partial_{x}u(y,t)\to\pm\infty$ as $y\to\pm 1$ given by Lemma 2.1 similarly implies that the function $(x,t)\mapsto t\,\phi(x,t):J_{1}\times(0,T]\mapsto\mathbb{R}$ extends continuously to $\overline{J_{1}}\times[0,T]$. Thus so does $\mathcal{H}:J_{1}\times(0,T]\mapsto(-\infty,\infty]$.
Next, we show that $\mathcal{H}$ restricted to the parabolic boundary of $J_{1}\times(0,T]$ is given by (2.12)–(2.13). The initial condition (2.12) and the first boundary condition of (2.13) are immediate, since $\phi(y,t)\to 0$ as $y\searrow-1$; we verify the second condition of (2.13). Denote by $v(y,t):=\partial_{x}u(y,t)$ the gradient of $u(t)$ at $y$. Let $p<q\in J_{1}$ and consider
(2.15) $A_{p,q}(t)=\int_{y\in(p,q)}u(y,t)\,\mathrm{d}y.$
@@
(2.16) $\partial_{t}A_{p,q}(s)$ $=\int_{y\in(p,q)}\partial_{t}u(y,s)\,\mathrm{d}y$
$=\int_{y\in(p,q)}\partial_{x}\!\left(\tan^{-1}(v)\right)\!(y,s)\,\mathrm{d}y$
$=\tan^{-1}(v(q,s)\!)-\tan^{-1}(v(p,s)\!)$
$=\phi(q,s)-\phi(p,s),$
@@
(2.18) $A_{p,q}(s)\to\mathcal{A}(q,s)\qquad\text{ and }\qquad\int_{s\in(0,t)}(\phi(q,s)-\phi(p,s)\!)\,\mathrm{d}s\to\int_{s\in(0,t)}\phi(q,s)\,\mathrm{d}s$
as $p\searrow-1$, the former for each $s\in[0,t]$, and hence that
(2.19) $\mathcal{A}(q,t)=\mathcal{A}(q,0)+\int_{s\in(0,t)}\phi(q,s)\,\mathrm{d}s.$
Taking $q\nearrow 1$, we have that
(2.20) $\mathcal{A}(q,t)\to\mathcal{A}(1,t)=\|u(t)\|_{\mathrm{L}^{1}(J_{1})}=\overline{\mathcal{A}}+\pi t$
@@
Finally, we show that $\mathcal{H}$ is smooth on $J_{1}\times(0,T)$. It suffices to show that $\partial_{t}\mathcal{A}$ and $\partial_{x}\mathcal{A}$ are smooth; but this is clear, since by the Dominated Convergence Theorem one can differentiate (2.19) and (2.8) to get that
(2.22) $\partial_{t}\mathcal{A}$ $=\phi$
@@
(2.23) $\partial_{x}\mathcal{A}$ $=u,$
which are both clearly smooth. ∎
@@
Proof.
@@
(2.26) $\partial^{2}_{x}\mathcal{A}$ $=v,$
(2.27) $\partial_{t}\phi$ $=\partial_{t}\!\left(\tan^{-1}(v)\!\right)=\frac{1}{1+v^{2}}\partial_{t}v$
@@
(2.28) $\partial_{x}^{2}\phi$ $=\partial_{x}^{2}\!\left(\tan^{-1}(v)\!\right)=\partial_{t}v.$
@@
(2.29) $\partial_{t}\mathcal{H}$ $=\partial_{t}\mathcal{A}-2\,\phi-2t\,\partial_{t}\phi=-\phi-\frac{2t}{1+v^{2}}\partial_{t}v$
@@
(2.30) $\partial^{2}_{x}\mathcal{H}$ $=\partial^{2}_{x}\mathcal{A}-2t\,\partial^{2}_{x}\phi=v-2t\,\partial_{t}v.$
@@
(2.31) $\partial_{t}\mathcal{H}-\frac{1}{1+v^{2}}\partial_{x}^{2}\mathcal{H}$ $=-\phi-\frac{v}{1+v^{2}}=-\tan^{-1}(v)-\frac{v}{1+v^{2}}-\frac{\pi}{2}\geq-\pi,$
@@
(2.32) ~$\widetilde{\mathcal{H}}(x,t)=\mathcal{H}(x,t)+\pi t$

### v2

@@ 1.2 Main results
Let $\nu=u_{0}\mathscr{L}^{1}+\nu_{\mathrm{sing}}\in\mathcal{M}_{*}(\mathbb{R})$ be a non-atomic real-valued Radon measure, decomposed into its absolutely continuous $\nu_{\mathrm{ac}}=:u_{0}\mathscr{L}^{1}$ and its singular $\nu_{\mathrm{sing}}$ part; here $\mathscr{L}^{1}$ is the Lebesgue measure on $\mathbb{R}$. Let $\Omega:=\mathbb{R}\setminus\operatorname{supp}(\nu_{\mathrm{sing}})$ and suppose that $u_{0}\in\mathrm{L}^{p}_{\mathrm{loc}}(\Omega)$ for some $1\leq p<\infty$ (recall that the weakest case of $p=1$ always holds by definition). Then there exists a Graphical Curve Shortening Flow $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(\mathbb{R}\times(0,\infty)\!)$ which observes the following modes of convergence as $t\searrow 0$:
$u(t)\mathscr{L}^{1}\rightharpoonup\nu\qquad$ $\text{ weakly in }\,\mathcal{M}(\mathbb{R})$ (1.15)
$u(t)\to u_{0}\qquad$ $\text{ strongly in }\,\mathrm{L}^{p}_{\mathrm{loc}}(\Omega).$ (1.16)
@@
    $u(t)\to u_{0}\qquad\textit{ strongly in }\,\mathrm{L}^{p}(\mathbb{R}).$ (1.17)
@@ 1.3 Acknowledgements
The author would like to thank his supervisor Prof. Peter M. Topping for the countless and invaluable discussions and suggestions which lead to this paper, and to thank Drs. Ben S. Lambert and Huy T. Nguyễn for a number of helpful comments. The author also thanks the anonymous referee at Discrete and Continuous Dynamical Systems for their corrections and suggestions, which greatly improved the exposition here.
@@ 2 A local Harnack quantity
Let $u_{0}\in\mathrm{C}^{0,1}(J_{1};\mathbb{R}_{\geq 0})$ be a positive Lipschitz function. We construct the Lipschitz curve $\gamma_{0}:\mathbb{R}\mapsto\mathbb{R}^{2}$ as illustrated in Fig. 2.1: for $x\in J_{1}$, set $\gamma_{0}(x)=(x,u_{0}(x)\!)$; for $x=1+y\in[1,\infty)$, set $\gamma_{0}(x)=(1,u_{0}(1)+y)$; and for $x=-1-y\in(-\infty,-1]$, set $\gamma_{0}(x)=(-1,u_{0}(-1)+y)$. Then the Chou–Zhu flow $0,\infty)\ni t\mapsto\gamma(t)\subset\mathbb{R}^{2}$ of $\gamma_{0}=:\gamma(0)$ admits a graphical representation for some $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,\infty)\!)$ with $u(t)\to u_{0}\in\mathrm{C}^{0}_{\mathrm{loc}}(J_{1})$ as $t\searrow 0$. Moreover, $u(y,t),\pm\partial_{x}u(y,t)\to\infty$ locally uniformly in $t\in(0,\infty)$ as $y\to\pm 1$.
@@
?figurename? 2.1: Construction of $\gamma_{0}:\mathbb{R}\mapsto\mathbb{R}^{2}$ from a given $u_{0}\in\mathrm{C}^{0,1}(J_{1};\mathbb{R}_{\geq 0})$.
?proofname? .
@@
  1. 1.
    First, we show that each $\gamma(t)$ for $t\in[0,T]$ is asymptotic to the vertical lines $V_{\pm 1}:=\left\{\pm 1\right\}\times\mathbb{R}\subset\mathbb{R}^{2}$. Applying the Avoidance Principle (Theorem A.7) to $\gamma$ and to huge shrinking circles placed either to the left of $V_{-1}$ or to the right of $V_{1}$ shows that $\gamma(t)$ must live within the strip $\overline{J_{1}}\times\mathbb{R}\subset\mathbb{R}^{2}$. Applying the Avoidance Principle again with huge circles placed below the xx–axis shows that $\gamma(t)$ must actually live within the half-strip $\overline{J_{1}}\times\mathbb{R}_{\geq 0}$. The Strong Maximum Principle implies that $\gamma(t)$ cannot touch the boundary of this half-strip.
@@
    $\mathfrak{o}(s):=\left\{(x,y)\in J_{1}\times\mathbb{R}\,\middle|\,\cosh\left(\frac{\pi y}{2}\right)=e^{-\pi^{2}s/4}\cos\left(\frac{\pi x}{2}\right)\right\}\subset\mathbb{R}^{2}.$ (2.2)
    Let $L>\|u_{0}\|_{\mathrm{L}^{\infty}(J_{1})}$; by translating $\mathfrak{o}(s)$ upwards to touch the line $H_{L}:=\mathbb{R}\times\{L\}\subset\mathbb{R}^{2}$ from above, and by considering the flow $0,|s|)\ni t\mapsto\mathfrak{o}(t+s)$, the Avoidance Principle ensures that the flow of $\gamma_{0}$ must lie outside the flow of the Angenent oval solution—see Fig. 2.2—for any $s<0$. In particular, a direct calculation (see e.g. [Sob24, §2.5.2]) shows that as $s\searrow$-$\infty$, the Angenent oval solutions approach the vertical translation of the Grim Reaper solution $0,\infty)\ni t\mapsto\mathfrak{g}(t)\subset\mathbb{R}^{2}$, given by
    $\mathfrak{g}(t):=\left\{(x,y)\in J_{1}\times\mathbb{R}\,\middle|\,e^{-\pi y/2}=e^{-\pi^{2}t/4}\cos\left(\frac{\pi x}{2}\right)\right\}\subset\mathbb{R}^{2},$ (2.3)
@@
    Next, we show that the curve $\gamma(t)$ is given by a graph in two senses. We first apply Angenent’s Intersection Principle (Theorem A.6) to $\gamma$ and to the static flow of the vertical lines $V_{r}:=\left\{r\right\}\times\mathbb{R}\subset\mathbb{R}^{2}$ for $r\in J_{1}$; see Fig. 2.2. There is initially a single intersection coming from the graphical representation $\gamma_{0}=\operatorname{graph}(u_{0})$, and for strictly positive times the curve $\gamma(t)$ must cross $V_{r}$ at least once; thus Angenent’s Intersection Principle ensures that the curve $\gamma(t)$ cross each $V_{r}$ exactly once. Hence $\gamma(t)=:\operatorname{graph}(u(t)\!)$ is a graph of a function $u(t):J_{1}\mapsto\mathbb{R}_{>0}$.
@@
    We second apply Angenent’s Intersection Principle to $\gamma$ and to the horizontal lines $H_{r^{\prime}}:=\mathbb{R}\times\left\{r^{\prime}\right\}\subset\mathbb{R}^{2}$; see Fig. 2.2. Let $\underline{r}^{\prime}\gg 1$ be such that $u(0,t)<\underline{r}^{\prime}$ for all $t\in[0,T]$. Then, similar to before, we have for all ${r^{\prime}}\geq\underline{r}^{\prime}$ that each $\gamma(t)$ for $t\in[0,T]$ intersects $H_{r^{\prime}}$ exactly twice, and therefore that the ends of $\gamma(t)$ are graphical over $\left\{\pm 1\right\}\times[\underline{r}^{\prime},\infty)\subset V_{\pm 1}$.
  3. 3.
@@
    Finally, we establish the convergence properties from regularity theory. We know from 2. that, say, for $t\in[0,T]$ the right-most end of $\gamma(t)$ can be written as a graph of $\widetilde{u}(t):[\underline{r}^{\prime},\infty)\mapsto\mathbb{R}$ over $\left\{1\right\}\times[\underline{r}^{\prime},\infty)\subset V_{1}$, which starts off as $\widetilde{u}(0)=0$ but which instantly lifts off from zero (i.e. $\widetilde{u}(t)>0$ for all $t\in(0,T]$). From 1., we know that the graphing function $\widetilde{u}(y,t)$ decays rapidly, say $|\widetilde{u}(y,t)|\leq\Psi(y)\to 0$, as $y\nearrow\infty$, since $\gamma(t)$ is squeezed between the vertical line $V_{1}$ and an upward-translated Grim Reaper (2.3).
@@
    Ecker & Huisken’s gradient estimate [EH91, 2.1 Theorem] implies that $\partial_{x}\widetilde{u}(y,t)$ remains uniformly bounded for, say, $(y,t)\in[\underline{r}^{\prime\prime},\infty)\times[0,T]$ for some $\underline{r}^{\prime\prime}\gg\underline{r}^{\prime}$. Standard parabolic regularity theory (see e.g. [LSU88, Kry96]) then ensures that the gradient $\partial_{x}\widetilde{u}(y,t)$ of the graphing function must also decay to zero as $y\nearrow\infty$ at the same rate as $\widetilde{u}(y,t)$: $|\partial_{x}\widetilde{u}(y,t)|\leq C\Psi(y)\to 0$. Written as a graph over $J_{1}\times\{0\}$, this implies that both $u(y,t)$ and $\partial_{x}u(y,t)$ blow up to infinity as $y\nearrow 1$, locally uniformly in $t\in(0,T]$. The argument for the left-most end of $\gamma(t)$ is similar. ∎
@@
$\gamma_{0}$$V_{r}$$H_{r^{\prime}}$
?figurename? 2.2: Left: Applying the Avoidance Principle (Theorem A.7) to the solid $\gamma$ and to a dashed Angenent oval solution $\mathfrak{o}$ in 1.
Right: Applying Angenent’s Intersection Principle (Theorem A.6) to the solid $\gamma$ and to a dashed vertical/horizontal line $V_{r}$$H_{r^{\prime}}$ in 2.
Remark 2.4 ($\mathrm{L}^{1}$–dominating function).
The proof of Lemma 2.1, in particular 1., shows the following: Let $L>\|u_{0}\|_{\mathrm{L}^{\infty}(J_{1})}$. Then the restricted solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,T])$ from Lemma 2.1 is dominated by the Grim Reaper solution (2.3) translated upwards by $L$; that is,
$u(x,t)\leq L+\frac{\pi t}{2}-\frac{2}{\pi}\log\left[\cos\left(\frac{\pi x}{2}\right)\right]\leq L+\frac{\pi T}{2}-\frac{2}{\pi}\log\left[\cos\left(\frac{\pi x}{2}\right)\right]$ (2.5)
Note that if one wished to apply the Avoidance Principle (Theorem A.7) directly to $\gamma$ and the Grim reaper solution, both of which are non-compact, then one would have first needed to establish the asymptotic behaviour of $\gamma$. Working with families of Angenent oval solutions, rather than directly with the Grim reaper solution, circumvents this technicality.
@@
For the purposes of 3. in Lemma 2.1, it was not important to clarify how the constant $C$ depended on the solution $\widetilde{u}$, as we only required the qualitative statement that $\partial_{x}\widetilde{u}(y,t)\to 0$ as $y\nearrow\infty$ locally uniformly in $t\in(0,\infty)$. In the proof of Theorem 1.14, it will be paramount to obtain quantitative estimates that depend on only weak information about the solution. Therefore, in §5.1 we will carry out a procedure similar to above, but with a little more care.
@@
Let $u_{0}\in\mathrm{C}^{0,1}(J_{1};\mathbb{R}_{\geq 0})$ be a positive Lipschitz function on $J_{1}$. The local GCSF starting from $u_{0}$ is the solution $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,\infty)\!)$ to the GCSF constructed in Lemma 2.1.
Note that a local GCSF $u$ extends to a function in $\mathrm{C}^{0}_{\mathrm{loc}}(J_{1}\times[0,\infty)\!)$, and is, by definition, positive.
@@
Let $u\in\mathrm{C}^{\infty}_{\mathrm{loc}}(J_{1}\times(0,\infty)\!)$ be a local GCSF. Define the area and angle [^3] The quantity $\phi(y,t)$ is the angle that the tangent to the curve $\gamma(t):=\operatorname{graph}(u(t)\!)$ at (y,u(y,t))(y,u(y,t)\!) makes with the downward vertical direction. functions $\mathcal{A}:J_{1}\times[0,\infty)\mapsto[0,\infty]$ and $\phi:J_{1}\times(0,\infty)\mapsto(0,\pi)$ by
$\mathcal{A}(y,t)$ $:=\int_{x\in(-1,y)}u(x,t)\,\mathrm{d}x$ (2.9)
@@
?figurename? 2.3: Illustration of $\mathcal{A}(y,t)$ and $\phi(y,t)$, where $\gamma(t):=\operatorname{graph}(u(t)\!)$.
The content of the following proposition is to confirm that $\mathcal{H}$ is smooth on $J_{1}\times(0,\infty)$, and extends continuously to $\overline{J_{1}}\times[0,\infty)$ with well-controlled boundary behaviour:
@@
$\mathcal{A}(-1,t)=0\qquad\text{ and }\qquad\mathcal{A}(1,t)=\|u(t)\|_{\mathrm{L}^{1}(J_{1})}.$ (2.15)
The boundedness of $\phi(y,t)$ and the locally uniform in $t\in(0,T]$–convergence of $\partial_{x}u(y,t)\to\pm\infty$ as $y\to\pm 1$ given by Lemma 2.1 similarly imply that the function $(x,t)\mapsto t\,\phi(x,t):J_{1}\times(0,T]\mapsto\mathbb{R}$ extends continuously to $\overline{J_{1}}\times[0,T]$. Thus so does $\mathcal{H}:J_{1}\times(0,T]\mapsto(-\infty,\infty]$.
Next, we show that $\mathcal{H}$ restricted to the parabolic boundary of $J_{1}\times(0,T]$, i.e. to ($J_{1}\times\left\{0\right\})\cup(\left\{\pm 1\right\}\times[0,T])$, is given by (2.13)–(2.14). The initial condition (2.13) and the first boundary condition of (2.14) are immediate, since $\phi(y,t)\to 0$ as $y\searrow-1$; we verify the second condition of (2.14). Denote by $v(y,t):=\partial_{x}u(y,t)$ the gradient of $u(t)$ at $y$. Let $(a,b)\Subset J_{1}$ and consider
$A_{a,b}(t):=\int_{y\in(a,b)}u(y,t)\,\mathrm{d}y.$
@@
$\partial_{t}A_{a,b}(s)$ $=\int_{y\in(a,b)}\partial_{t}u(y,s)\,\mathrm{d}y$ (2.16)
$=\int_{y\in(a,b)}\partial_{x}[\tan^{-1}(v)](y,s)\,\mathrm{d}y$
$=[\tan^{-1}(v)](b,s)-[\tan^{-1}(v)](a,s)$
$=\phi(b,s)-\phi(a,s),$
@@
$A_{a,b}(s)\to\mathcal{A}(b,s)\qquad\text{ and }\qquad\int_{s\in(0,t)}[\phi(b,s)-\phi(a,s)]\,\mathrm{d}s\to\int_{s\in(0,t)}\phi(b,s)\,\mathrm{d}s$ (2.18)
as $a\searrow-1$, the former for each $s\in[0,t]$, and hence that
$\mathcal{A}(b,t)=\mathcal{A}(b,0)+\int_{s\in(0,t)}\phi(b,s)\,\mathrm{d}s.$ (2.19)
Taking $b\nearrow 1$, we have that
$\mathcal{A}(b,t)\to\mathcal{A}(1,t)=\|u(t)\|_{\mathrm{L}^{1}(J_{1})}=\overline{\mathcal{A}}+\pi t$ (2.20)
@@
Finally, we show that $\mathcal{H}$ is smooth on $J_{1}\times(0,T)$. It suffices to show that both $\partial_{t}\mathcal{A}$ and $\partial_{x}\mathcal{A}$ are smooth; but this is clear, since by the Dominated Convergence Theorem one can differentiate (2.19) and (2.9) to get that
$\partial_{t}\mathcal{A}$ $=\phi$ (2.22)
@@
$\partial_{x}\mathcal{A}$ $=u$ (2.23)
respectively, which are both clearly smooth. ∎
@@
Our justification for calling $\mathcal{H}$ a “Harnack quantity”, and consequently (2.25) a “Harnack inequality”, both for the $\mathcal{H}$ here and in [ST24a], is in its links to other Harnack quantities for other geometric flows—see, for example, Remark 2.34.
For a concrete link between Hamilton’s Harnack Inequality [Ham93, Main Theorem A] for the MCF and an adapted version of our Harnack inequality [ST24a, (2.11)]/(2.25), the interested reader is directed to [ST26, §5], where Topping & the author show, in the setting of the CSF of convex curves, that the two Harnack inequalities imply a common pointwise curvature estimate.
Proof of Theorem 2.24.
@@
$\partial^{2}_{x}\mathcal{A}$ $=v,$ (2.26)
$\partial_{t}\phi$ $=\partial_{t}[\tan^{-1}(v)]=\frac{1}{1+v^{2}}\partial_{t}v$ (2.27)
@@
$\partial_{x}^{2}\phi$ $=\partial_{x}^{2}[\tan^{-1}(v)]=\partial_{t}v.$ (2.28)
@@
$\partial_{t}\mathcal{H}$ $=\partial_{t}\mathcal{A}-2\,\phi-2t\,\partial_{t}\phi=-\phi-\frac{2t}{1+v^{2}}\partial_{t}v$ (2.29)
@@
$\partial^{2}_{x}\mathcal{H}$ $=\partial^{2}_{x}\mathcal{A}-2t\,\partial^{2}_{x}\phi=v-2t\,\partial_{t}v.$ (2.30)
@@
$\partial_{t}\mathcal{H}-\frac{1}{1+v^{2}}\partial_{x}^{2}\mathcal{H}$ $=-\phi-\frac{v}{1+v^{2}}=-\tan^{-1}(v)-\frac{v}{1+v^{2}}-\frac{\pi}{2}\geq-\pi,$ (2.31)
@@
$\widetilde{\mathcal{H}}(x,t)=\mathcal{H}(x,t)+\pi t$ (2.32)